OpenStudy (anonymous):

Find an equation of a parabola with a vertex at the origin and directrix y = –1.5

OpenStudy (mertsj):

Since the directrix is y = -1.5 we know that the focus is at (0,1.5) and the value of p is 1.5. So we can plug into the equation x^2=4py to find the equation.

OpenStudy (anonymous):

is the answer x=1/6y^2

OpenStudy (anonymous):

i don't understand how to do it??

Directrix (directrix):

General Equation: (x – h)² = 4p(y – k) where (h,k) is the vertex of the parabola and p is the absolute value of the distance from the vertex to the directrix. So, p = 3/2. (x – h)² = 4p(y – k) (x - 0)² = 4*(3/2)*(y - 0) x² = 6*y y = (1/6) * x² @nsschwab23 A good study resource is at this link: http://tinyurl.com/cgdfblf